On the number line x=1/4 and y=11/12. The point z divides the segment from x to y into two parts such that the distance from x to z is 3/8 of the distance from z to y. Find the distance from z to y.
Answer by greenestamps(13200)



If the ratio of XZ to ZY is 3:8, then XZ is 3/11 of the distance XY and ZY is 8/11 of XY.

Then, since XY is 11/12-1/4 = 11/12-3/12 = 8/12 = 2/3, the distance from Z to Y is
(8/11)(2/3) = 16/33.